61500 64100 72500 92600 86700 57600 74500 79100 68200 64100 79900 77200 65200 77700 53300 70200 72100 80200 49700 67800 49100 55500 82400 52400 Construct a grouped frequency distribution table (GFDT) for this data set. Your classes should be labeled using interval notation. Hint: The first Data Range is: [45000,49900]. The next bin should start with the next number. Data range Frequency
Answer #1

Steps to Create Frequency Distribution:

1) Width of the class is given to be (49900-45000 = 4900).

2) Find the data range by subtracting the minimum range value from the maximum data value. In this case, the data range is 92600-45000 = 47600

3) Find the desired number of groups by dividing the data range by class width. Ie 47600/4900 = 9.7. Ceiling of this value is 10 and we will create 10 groups each of 4900 width starting from 45001.

4) Then we count the no of data set in each interval and generate the frequency table.

Range 45001-49900 49901-54800 54801-59700 59701-64600 64601-69500 69501-74400 74401-79300 79301-84200 84201-89100 89101-94000 Total 2 2 4 4 1 1 28